alones (alones):

Write the quadratic equation in factored form. Be sure to write the entire equation.

alones (alones):

\[x^2 + x - 12 = 0\]

alones (alones):

Liek this (x−3)(x+4)=0 ?

OpenStudy (roxy.girl):

x^2+x-12=0 x^2+4x-3x-12 x(x+4)-3(x+4) (x-3)(x+4)

alones (alones):

thansk roxy<3

OpenStudy (roxy.girl):

You welcome <3 o.o It isn't showing I am typing :(